The diameter of a Ferris wheel is 10 meters. From a platform 2 meters above the ground, you board the Ferris wheel at the bottom. Determine an equation for a rider's height in relation to time if it takes 35 seconds to reach the summit. Sketch the graph.

1
Expert's answer
2022-02-22T16:06:43-0500

Solution:

"d=10\ m
\\ r=5\ m"

So, amplitude"=a=5"

It takes 35 seconds to reach the summit.

So, 70 seconds for one complete revolution.

"b=\dfrac{2\pi}{70}=\dfrac{\pi}{35}"

The lowest point is 2 m off the ground and the radius is 5 m, so the midline will be at 7 m.

Thus, equation is: "h(t)=-5\cos\dfrac{\pi}{35}t+7"

Assume it as "y=-5\cos\dfrac{\pi}{35}x+7"

Its graph is: